OpenStudy (anonymous):

Air is 21% oxygen by volume. Oxygen has a density of 1.31 g/L. What is the volume, in Liters, of a room that holds enough air to contain 55 kg of oxygen?

OpenStudy (kc_kennylau):

(a) How many liters is 55 kg of oxygen? (b) How many liters of air contain that many liters of oxygen?

OpenStudy (anonymous):

Also there was the hint.. If air is 21% oxygen by volume, 1L air= 0.21L oxygen

OpenStudy (anonymous):

I think my teacher wanted me to use the conversion factor or the dimensional analysis or whatever but.. idk where to start

OpenStudy (kc_kennylau):

so 1.31g of oxygen consumes 1L of space.

OpenStudy (kc_kennylau):

how much space would 55kg of oxygen consume?

OpenStudy (anonymous):

idk...

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@myininaya

OpenStudy (wolf1728):

1/.21 = 4.7619047619 So a room that contains 55 kg of oxygen, contains 55*4.7619047619 = 261.9 kg of air. Assuming air is same density as air (and it ISN'T) then 261.9kg of air = 261,900 grams of air = 261,900 liters of air

OpenStudy (anonymous):

You sure?

OpenStudy (anonymous):



OpenStudy (anonymous):

I kinda did this

OpenStudy (wolf1728):

=55,000 / (1.31*.21) =199927.29916394

OpenStudy (anonymous):

I mean the room will be filled with air... so finding the amount of oxygen and proportioning it to the volume of the room, interchangeably, the room, would tell me this riiight?

OpenStudy (wolf1728):

55 kg oxygen x (1,000 g / kg) x (1 liter / 1.31 g) x 4.7619 l air / l oxygen